Let A=i+3j-2k

and
B=4i-2j+4k
, find
(2A+B)(A-2B)

Answer by Alan3354(69443)

Let A=i+3j-2k

and
B=4i-2j+4k
, find
(2A+B)(A-2B)
--------------
2A + B = (2i+6j-4k) + 4i-2j+4k = 6i+4j
=============================
A - 2B = i+3j-2k - (8i-4j+8k) = -7i+7j-10k
---------------------------
Do you mean the cross product? Or the dot product?